# Spin Theory Based on the Extended Least Action Principle and Information Metrics: Quantization, Entanglement, and Bell Test With Time Delay

### v2

@@ VII.2 Limitations
There are several limitations we need to point out for future investigations. The first challenge is to explain why the magnitude of intrinsic angular momentum of an electron is $\hbar/2$. Although Appendix B gives a reasonable derivation, which is also based on the extended least action principle, there is a strong assumption on random circular motion due to vacuum fluctuation. A more intuitive model with weaker assumptions is desirable.
@@
Secondly, assumption 4 suggests that when the interaction between the electron and an inhomogeneous magnetic field in a Stern-Gerlach apparatus takes place, the order of relative entropy monotonically increases, so that the direction of intrinsic angular momentum tends to align with the direction of the magnetic field. This assumption appears quite arbitrary, but can be intuitively explained as follows. The key is not to consider the electron as an idealized point particle. Instead it has non-zero size and is distributed with a spatial volume. We do not assume it is a rigid rotating body. Instead, it can be divided into many small mass elements $\delta m$. Each mass element is also a charge element $\delta e$, and experiences Larmor precession. Note that in Larmor precession, the angular frequency is $\omega=\delta eB_{z}/\delta m$. In a homogeneous magnetic field, $\omega$ is a constant for each charge element and independent of the angle $\theta$. In this case, the probability density function $p_{m}(\theta)$ should remain unchanged for the whole electron. But in an inhomogeneous field $B_{z}$, because the electron has spatial size, each charge element will experience a different magnitude of magnetic field. Consequently, different parts of the electron will have different angular frequencies with difference of $\Delta\omega$. The upper parts of the electron along the direction of $B_{z}$ tend to precess faster than the lower parts of the electron. Effectively, the orientation of overall intrinsic angular momentum is “pulled” closer to align with the $z$-axis. This corresponds to choosing a larger value of $m$. The spin theory based on stochastic mechanics Beyer seems to give similar results.
@@
To estimate $\Delta\omega$, consider that in a typical Stern-Gerlach experiment, the gradient of magnetic field is of the order of $10^{3}$ Tesla/m. The size of electron is of order of $10^{-18}$m. Thus, the difference of magnetic field experienced at the upper part and lower part of the electron is about $\Delta B_{z}=10^{-15}$ Tesla. Then, the the Larmor precession frequency difference $\Delta\omega$ is at the order of $10^{-3}$Hz. On the other hand, typical Larmor frequency $\omega$ for an electron is $10^{10}$Hz. Thus, $\Delta\omega/\omega=10^{-13}$. This is an extremely small ratio. It raises the question on whether such a small ratio can cause the orientation of overall intrinsic angular momentum to tilt closer to align with the $z$-axis. A more rigorous theory to confirm this assumption is needed. Either considering the electron as an idealized charged point particle, or as a charge body with finite spatial size, seems insufficient as an accurate physical model for the electron. We speculate that using field theory could be a better framework to describe the physical model of electron, which is beyond the scope of this work but has been studied in Sebens .
@@ VII.3 Conclusions
From the extended least action principle, and with additional assumptions on the intrinsic angular momentum of an electron, particularly its random orientations, we are able to develop a theory that recovers the quantum formulation for electron spin. The key component of the theoretical framework here is the introduction of relative entropy for both translational and rotational random fluctuations. The two-level quantization of spin measurement outcomes is a mathematical consequence based on Assumption 4. We also obtain the same formulation as standard quantum mechanics when the direction of magnetic field of a second Stern-Gerlach apparatus is rotated with an angle from the direction of magnetic field of the first Stern-Gerlach apparatus. Recursively applying the extended least action principle, we have derived the Schrödinger-Pauli equation.
@@
An important result is that we provide an intuitive physical model and formulation to explain the entanglement phenomenon between two electron spins. The root cause of entanglement is the correlation between the orientations of the intrinsic angular momenta of the two electrons. Essentially, the correlation is established between two random variables due to previous interactions. Since the orientation of the angular momentum is an intrinsic local property of the electron, the correlation can be maintained even though the two electrons are space-like separated and have no further interaction. Mathematically, we give an equivalent formulation of the four Bell states using probability density function instead of wave function. Using the probability density functions for a pair of entangled electrons, we prove that the Bell-CHSH inequality is violated. The violation of Bell-CHSH inequality is due to the same root cause of entanglement, that is, correlation of the orientations of the angular momenta which is preserved even when the two electrons are separated.
@@
When certain conditions are not taken to the limits, the present theory predicts results that are different from standard quantum mechanics. For instance, if the Stern-Gerlach apparatus is set up with a sufficiently weak gradient of the magnetic field along the $z$-axis, we expect the electron detector screen will exhibit continuous distribution of displacements along the $z$-axis, instead of only two discrete lines. Another more interesting experiment proposed is to add a delay before Bob’s measurement in the typical Bell test experiment, which could result in the non-violation of the Bell-CHSH inequality. The second experiment shows the effect of time dependency in measurements involving spin entanglement. If confirmed by experiment, such an effect can impose a limitation of the application of spin entanglement in quantum computing and quantum information.
@@
The interplay of the extended least action principle and the information metrics for vacuum fluctuations prove to be very fruitful. We have shown the theoretical framework based on the extended least action principle has successfully recovered the non-relativistic quantum mechanics Yang2023 , the relativistic quantum scalar field theory Yang2024 , and in the paper, the quantum theory for electron spin. We believe the results in this paper are more interesting and because we can explain the entanglement phenomenon with an intuitive physical picture, and propose experiment to test potential new results different from standard quantum mechanics.
@@
There are clear limitations in the present theory with the assumptions on the detailed physical model of electron spin. A more rigorous theory must be developed to further justify these assumptions. This paper is intended to be an incremental step for the investigation of spin theory using the information metrics for vacuum fluctuations.
@@ Appendix A Deriving Basic Quantum Formulation From the Extended Least Action Principle
Basic quantum formulation can be derived by recursively applying the extended least action principle in two steps. First, we consider the dynamics of a system with an infinitesimal time internal $\Delta t$ due to vacuum fluctuation only. Define the probability for the system to transition from a 3-dimensional space position $\mathbf{x}$ to another position $\mathbf{x}+\mathbf{w}$, where $\mathbf{w}=\Delta\mathbf{x}$ is the displacement in 3-dimensional space due to fluctuations, as $\wp(\mathbf{x}+\mathbf{w}|\mathbf{x})d^{3}\mathbf{w}$. The expectation value of classical action is $S_{c}=\int\wp(\mathbf{x}+\mathbf{w}|\mathbf{x})Ld^{3}\mathbf{w}dt$. Since we only consider the vacuum fluctuations, the Lagrangian $L$ only contains the kinetic energy, $L=\frac{1}{2}m\mathbf{v}\cdot\mathbf{v}$. For an infinitesimal time internal $\Delta t$, one can approximate the velocity $\mathbf{v}=\mathbf{w}/\Delta t$. This gives
@@
Suppose the initial condition is unknown, and define $\rho({x},t)$ as the probability density for finding a particle in a given volume of the configuration space. The probability density must satisfy the normalization condition $\int\rho(\mathbf{x},t)d^{3}\mathbf{x}=1$, and the continuity equation
@@
through fixed point variation with respect to $\rho$ and $S$, respectively. Note that $A_{c}$ and $S$ are different physical variables, where $A_{c}$ can be considered as the ensemble average of classical action and $S$ is a generation function that satisfied $\mathbf{p}=\nabla S$ Yang2023 .
@@
Eq. (84) contains the term related to Fisher information for the probability density FriedenBook but bears much more physical significance than Fisher information. Inserting (81) and (84) into (1), and performing the variation procedure on $I$ with respect to $S$ gives the continuity equation, while variation with respect to $\rho$ leads to the quantum Hamilton-Jacobi equation,
@@
The last term in (85) is the Bohm quantum potential Bohm1952 . The Bohm potential is considered responsible for the non-locality phenomenon in quantum mechanics Bohm2 . Historically, its origin is mysterious. Here we show that it originates from the information metrics related to relative entropy, $I_{f}$.
As pointed out in Yang2023 , the choice of relative entropy for $I_{f}$ can be Renyi divergence or Tsallis divergence, which results in a family of Schrödinger equations that depends on the order of relative entropy $\alpha$. When $\alpha=1$, the regular Schrödinger equation is recovered. The flexibility to choose general relative entropy definitions is a very useful mathematical tool for further theoretical investigation of more advanced quantum theory, as we have shown in the present work.
@@ Appendix B Intrinsic Angular Momentum Due to Vacuum Fluctuations
Assumption 1 on vacuum fluctuations in Section II does not specify the details of the fluctuating motion. The electron exhibits random displacement due to vacuum fluctuation and such displacement is described as a vector. The displacement vector is intrinsically local in the sense that it is independent of the reference of origin of the global coordinate system. Given a local point and a infinitesimal time period, the displacement vector can be just a translational motion, i.e., the direction of velocity is in parallel with the displacement vector. However, there is no reason that the direction of velocity must be in parallel with the displacement vector. When the velocity vector has a component perpendicular to the displacement vector, the electron will exhibit rotational movement that can be characterized by an angular momentum. Here we will further assume that rotational movement is circular. Thus, the vacuum fluctuation not only causes the translational displacement, but also comprises components of motion as circular rotations. Essentially, the vacuum fluctuation produces virtual circular rotations with random radius. Such a model has been proposed in Niehaus16 ; Niehaus22 , but no derivation was provided on the magnitude of intrinsic angular momentum. Here, with the help of the extended least action principle, we will show that such random circular rotations give rise to the intrinsic angular momentum with magnitude of $\hbar/2$.
Denote the radius vector of circular motion in an infinitesimal time period $\Delta t$ as $\vec{u}$ and the velocity vector as $\dot{\vec{u}}$. Here we will only consider the circular motion of the electron such that $\dot{\vec{u}}$ is perpendicular to $\vec{u}$. The angular momentum is the cross product of $\vec{u}$ and $m\dot{\vec{u}}$
@@
The orientation of $\vec{L}_{s}$ is completely random. We choose the coordinate plane perpendicular to $\vec{L}_{s}$ for further analysis of the circular motion. Here, the radius $u$ is a random variable. Let $\omega$ be the angular frequency. Then, the magnitude of velocity is $\dot{u}=\omega u$, and the magnitude of angular momentum is $L_{s}=m\omega u^{2}$. Since the radius $u$ is a random variable, we denote $p(u)$ as the probability density for radius being $u$. It must satisfy the normalization condition
@@
For this circular motion, the Lagrangian $L$ only contains the kinetic energy, $L=\frac{1}{2}m(\omega u)^{2}$. For an infinitesimal period of time $\Delta t$, the expectation value of classical action is
@@
Note that $d\varphi=\omega dt$ where $\varphi$ is the angle of circular rotation. In the infinitesimal period of time $\Delta t$, the electron rotates $\Delta\varphi=\omega\Delta t$. We can rewrite (89) as
@@
Thus, by assuming that the vacuum fluctuations cause random circular motions for the electron, we show that the electron possesses an intrinsic angular momentum with magnitude of $\hbar/2$. Such intrinsic angular momentum is local in the sense that it is independent of the global orbital movement. Its orientation is completely random. These properties have been summarized as Assumption 3 in Section III. But based on the derivation showing in this appendix, we can replace Assumption 3 with
@@
Appendix C Proof of (36) in Standard Quantum Mechanics

### v3

@@ VII.2 Limitations
There are several limitations that we need to point out for future investigations. The first challenge is to explain why the magnitude of intrinsic angular momentum of an electron is $\hbar/2$. Although Appendix B gives a reasonable derivation, which is also based on the extended least action principle, there is a strong assumption on random circular motion due to vacuum fluctuation. A more intuitive model with weaker assumptions is desirable.
@@
Secondly, Assumption 4 suggests that when the interaction between the electron and an inhomogeneous magnetic field takes place in a Stern-Gerlach apparatus, the order of relative entropy monotonically increases, so that the direction of intrinsic angular momentum tends to align with the direction of the magnetic field. This assumption seems quite arbitrary. Here, we give a possible intuitive explanation. The key is not to consider the electron as an idealized point particle. Instead, it has non-zero size and is distributed with a spatial volume. We do not assume that it is a rigid rotating body. Instead, it can be divided into many small mass elements $\delta m$. Each mass element is also a charge element $\delta e$, and experiences Larmor precession. The Larmor precession frequency is $\omega=\delta eB_{z}/\delta m$. In a homogeneous magnetic field, $\omega$ is a constant for each charge element and independent of angle $\theta$. In this case, the probability density function $p_{m}(\theta)$ should remain unchanged for the entire electron. But in an inhomogeneous field $B_{z}$, because the electron has spatial size, each charge element will experience a different magnitude of the magnetic field. Consequently, different parts of the electron will have different precession frequencies with $\Delta\omega\propto\Delta B_{z}$. The upper parts of the electron along the direction of $B_{z}$ tend to precess faster than the lower parts of the electron. Effectively, the orientation of the overall intrinsic angular momentum is “pulled” closer to align with the $z$-axis. This corresponds to choosing a larger value of $m$. Spin theory based on stochastic mechanics Beyer seems to give similar results.
@@
To estimate $\Delta\omega$, consider that in a typical Stern-Gerlach experiment, the gradient of magnetic field is of the order of $10^{3}$ Tesla/m. The size of the electron is of the order of $10^{-18}$ m. Thus, the difference in magnetic field experienced at the upper part and the lower part of the electron is about $\Delta B_{z}=10^{-15}$ Tesla. Then, the Larmor precession frequency difference $\Delta\omega$ is on the order of $10^{-3}$Hz. On the other hand, the typical Larmor frequency $\omega$ for an electron is $10^{10}$Hz. Thus, $\Delta\omega/\omega=10^{-13}$. This is an extremely small ratio. It raises the question whether such a small ratio can cause the orientation of the overall intrinsic angular momentum to tilt closer to align with the $z$-axis. A more rigorous theory is needed to confirm this assumption. Either considering the electron as an idealized charged point particle or as a charge body with finite spatial size, seems insufficient as an accurate physical model for the electron. We speculate that using field theory could be a better framework to describe the physical model of an electron, which is beyond the scope of this work but has been studied in Sebens .
@@ VII.3 Conclusions
From the extended least action principle, and with additional assumptions on the intrinsic angular momentum of an electron, particularly its random orientations, we are able to develop a theory that recovers the quantum properties for electron spin. The key component of the theoretical framework here is the introduction of relative entropy for both translational and rotational random fluctuations. The two-level quantization of spin measurement results is a mathematical consequence when the order of relative entropy approaches a limit. We also obtain the same formulation as standard quantum mechanics when the direction of magnetic field of a second Stern-Gerlach apparatus is rotated with an angle from the direction of magnetic field of the first Stern-Gerlach apparatus. Recursively applying the extended least action principle, we have derived the Schrödinger-Pauli equation.
@@
An important result here is that we provide an intuitive physical model and formulation to explain the entanglement phenomenon between two electron spins. The root cause of entanglement is the correlation between the orientations of the intrinsic angular momenta of the two electrons. The correlation is established between two random variables due to previous interactions. Since the orientation of the angular momentum is an intrinsic local property of the electron, the correlation can be maintained even though the two electrons are space-like separated and have no further interaction. Mathematically, we give an equivalent formulation of the four Bell states using probability density function instead of wave function. Using the probability density functions for a pair of entangled electrons, we prove that the Bell-CHSH inequality is violated. The violation of Bell-CHSH inequality is due to the same root cause of entanglement, that is, correlation of the orientations of the angular momenta which is preserved even when the two electrons are separated.
@@
When certain conditions are not taken to the limits, the present theory gives certain results that are different from standard quantum mechanics. For instance, if the Stern-Gerlach apparatus is set up with a sufficiently weak gradient of the magnetic field along the $z$-axis, we expect the electron detector screen to exhibit a continuous distribution of displacements along the $z$-axis, instead of only two discrete lines. Another more interesting experiment proposed is to add a delay before Bob’s measurement in the typical Bell test experiment, which could result in the non-violation of the Bell-CHSH inequality. The second experiment shows the effect of time dependency of the measurement results involving spin entanglement. If confirmed by experiment, such an effect can impose a limitation on the application of spin entanglement in quantum computing and quantum information.
@@
The interplay of the extended least action principle and the information metrics for vacuum fluctuations is proved to be a valuable theoretical framework. We have shown that the theoretical framework based on the extended least action principle has successfully recovered non-relativistic quantum mechanics Yang2023 , the relativistic quantum scalar field theory Yang2024 , and in this paper the quantum theory for electron spin. We believe that the results in this paper are interesting because we can explain the entanglement phenomenon with an intuitive physical model, and the spin-entanglement model implies potential new results that are falsifiable with the modified Bell test with time delay.
@@
There are clear limitations in the present theory with the assumptions on the detailed physical model of electron spin. A more rigorous theory must be developed to further justify these assumptions. This paper is intended to be an intermediate step in the investigation of spin theory using information metrics of vacuum fluctuations.
Acknowledgements.
The author thanks the anonymous referees for their valuable comments, which motivate the author to clarify the choice of the Tsallis divergence and the difference between classical correlation and spin entanglement.
@@ Appendix A Deriving Basic Quantum Formulation From the Extended Least Action Principle
Basic quantum formulation can be derived by recursively applying the extended least action principle in two steps. First, we consider the dynamics of a system with an infinitesimal time internal $\Delta t$ due to only vacuum fluctuation. Define the probability that the system will transition from a 3-dimensional space position $\mathbf{x}$ to another position $\mathbf{x}+\mathbf{w}$, where $\mathbf{w}=\Delta\mathbf{x}$ is the displacement in a 3-dimensional space due to fluctuations, as $\wp(\mathbf{x}+\mathbf{w}|\mathbf{x})d^{3}\mathbf{w}$. The expectation value of the classical action is $S_{c}=\int\wp(\mathbf{x}+\mathbf{w}|\mathbf{x})Ld^{3}\mathbf{w}dt$. Since we only consider the vacuum fluctuations, the Lagrangian $L$ only contains the kinetic energy, $L=\frac{1}{2}m\mathbf{v}\cdot\mathbf{v}$. For an infinitesimal time internal $\Delta t$, one can approximate the velocity $\mathbf{v}=\mathbf{w}/\Delta t$. This gives
@@
Suppose that the initial condition is unknown and define $\rho({x},t)$ as the probability density for finding a particle in a given volume of the configuration space. The probability density must satisfy the normalization condition $\int\rho(\mathbf{x},t)d^{3}\mathbf{x}=1$, and the continuity equation
@@
through fixed point variation with respect to $\rho$ and $S$, respectively. Note that $A_{c}$ and $S$ are different physical variables, where $A_{c}$ can be considered as the ensemble average of classical action and $S$ is a generation function that satisfies $\mathbf{p}=\nabla S$ Yang2023 .
@@
Eq. (84) contains the term related to Fisher information for the probability density Frieden but is much more physical significant than Fisher information. Inserting (81) and (84) into (1), and performing the variation procedure on $I$ with respect to $S$ gives the continuity equation, while variation with respect to $\rho$ leads to the quantum Hamilton-Jacobi equation,
@@
The last term in (85) is the Bohm quantum potential Bohm1952 . The Bohm potential is considered responsible for the non-locality phenomenon in quantum mechanics Bohm2 . Historically, its origin is mysterious. Here we show that it originates from information metrics related to relative entropy, $I_{f}$.
As noted in Yang2023 , the choice of relative entropy for $I_{f}$ can be Renyi divergence or Tsallis divergence, which results in a family of Schrödinger equations that depends on the order of relative entropy $\alpha$. When $\alpha=1$, the regular Schrödinger equation is recovered. The flexibility to choose general relative entropy definitions is a very useful mathematical tool for further theoretical investigation of more advanced quantum theory, as we have shown in the present work.
@@ Appendix B Intrinsic Angular Momentum Due to Vacuum Fluctuations
Assumption 1 on vacuum fluctuations in Section II does not specify the details of the fluctuating motion. The electron exhibits random displacement due to vacuum fluctuation, and such displacement is described as a vector. The displacement vector is intrinsically local in the sense that it is independent of the reference of origin of the global coordinate system. Given a local point and an infinitesimal time period, the displacement vector can be just a translational motion, i.e. the direction of velocity is in parallel with the displacement vector. However, there is no reason that the direction of velocity must be in parallel with the displacement vector. When the velocity vector has a component perpendicular to the displacement vector, the electron will exhibit rotational movement that can be characterized by an angular momentum. Here we will further assume that rotational movement is circular. Thus, the vacuum fluctuation not only causes the translational displacement, but also comprises components of motion such as circular rotations. Essentially, the vacuum fluctuation produces virtual circular rotations with random radius. This model has been proposed in Niehaus16 ; Niehaus22 , but no derivation of the magnitude of the intrinsic angular momentum has been provided. Here, with the help of the extended least action principle, we will show that such random circular rotations give rise to the intrinsic angular momentum of magnitude $\hbar/2$.
Denote the radius vector of the circular motion in an infinitesimal time period $\Delta t$ as $\vec{u}$ and the velocity vector as $\dot{\vec{u}}$. Here we will only consider the circular motion of the electron such that $\dot{\vec{u}}$ is perpendicular to $\vec{u}$. The angular momentum is the cross product of $\vec{u}$ and $m\dot{\vec{u}}$
@@
The orientation of $\vec{L}_{s}$ is completely random. We choose the coordinate plane perpendicular to $\vec{L}_{s}$ for a further detailed analysis of the circular motion. Here, the radius $u$ is a random variable. Let $\omega$ be the angular frequency. Then, the velocity magnitude is $\dot{u}=\omega u$, and the angular momentum magnitude is $L_{s}=m\omega u^{2}$. Since the radius $u$ is a random variable, we denote $p(u)$ as the probability density of the radius $u$. It must satisfy the normalization condition
@@
For this circular motion, the Lagrangian $L$ contains only the kinetic energy $L=\frac{1}{2}m(\omega u)^{2}$. For an infinitesimal period of time $\Delta t$, the expectation value of classical action is
@@
Note that $d\varphi=\omega dt$ where $\varphi$ is the angle of circular rotation. In an infinitesimal period of time $\Delta t$, the electron rotates $\Delta\varphi=\omega\Delta t$. We can rewrite (89) as
@@
Thus, by assuming that vacuum fluctuations cause random circular motions for the electron, we show that the electron possesses an intrinsic angular momentum with magnitude of $\hbar/2$. Such intrinsic angular momentum is local in the sense that it is independent of the global orbital movement. Its orientation is completely random. These properties have been summarized as Assumption 3 in Section III. But based on the derivation showing in this appendix, we can replace Assumption 3 with
@@
Appendix C The Choice of Relative Entropy
Eq. (8) is defined with the Tsallis divergence, a one-parameter generalization of the Kullback-Leilber divergence. Alternatively, we can define $I_{f}$ with another popular one-parameter generalization of the Kullback-Leilber divergence, the Rényi divergence. Here we show that defining $I_{f}$ with the Rényi divergence leads to the same results in spin quantization. The Rényi divergence is defined as
$I^{R}_{f}=\frac{1}{\alpha-1}ln(\int_{0}^{\pi}\int_{0}^{\Delta\varphi}\frac{p^{\alpha}(\theta)}{\sigma^{\alpha-1}}d\theta d\varphi).$ (97)
The total action in (9) becomes
$A_{t}=-\frac{1}{2}g_{s}L_{s}\int p(\theta)\cos(\theta)d\theta d\varphi+\frac{\hbar}{2(\alpha-1)}ln(\int\frac{p^{\alpha}(\theta)}{\sigma^{\alpha-1}}d\theta d\varphi).$ (98)
$\frac{1}{2}g_{s}L_{s}\cos(\theta)+\frac{\alpha\hbar}{2(\alpha-1)F_{\alpha}[p]}[\frac{p(\theta)}{\sigma}]^{\alpha-1}=0,$ (99)
where $F_{\alpha}[p]$ is a functional that depends on the functional form of $p$ but not on the parameter $\theta$ and $\varphi$,
$F_{\alpha}[p]=\int\frac{p^{\alpha}(\theta)}{\sigma^{\alpha-1}}d\theta d\varphi.$ (100)
Given the normalization condition, one recognize that
$\lim_{\alpha\to 1}F_{\alpha}[p]=\int p\theta d\varphi=1,$ (101)
which is independent of the functional form of $p$. From (99), we obtain
$p(\theta)=\sigma\{\frac{(\alpha-1)g_{s}L_{s}}{\alpha\hbar}F_{\alpha}[p]\cos(\theta)\}^{\frac{1}{\alpha-1}}=\frac{1}{Z_{\alpha}}\{F_{\alpha}[p]\cos(\theta)\}^{\frac{1}{\alpha-1}},$ (102)
Using the definition (12), the probability density (102) is rewritten as
$p_{m}(\theta)=\frac{1}{Z_{m}}(F_{m}[p]\cos(\theta))^{2m},m\in\mathbb{N}.$ (103)
Since $\lim_{m\to\infty}F_{m}[p]=1$, which is independent of the functional form of $p$, we still have
which is the same as (15). It shows that only two discrete measurement results can be obtained.
To conclude the discussion on the choice of relative entropy, suppose that we choose the K-L divergence, the total action becomes
$A_{t}=-\frac{1}{2}g_{s}L_{s}\int p(\theta)\cos(\theta)d\theta d\varphi+\frac{\hbar}{2}\int p(\theta)ln\frac{p(\theta)}{\sigma}d\theta d\varphi.$ (105)
Performing the variation procedure, we eventually obtain
$p(\theta)=\frac{1}{Z}e^{\frac{g_{s}L_{s}}{\hbar}\cos(\theta)}.$ (106)
This form of $p(\theta)$ does not lead to the conclusion of discrete measurement results. Although the K-L divergence is the limit of the Tsallis divergence or the Rényi divergence when $m\to\infty$ (or $\alpha\to 1$), note that in obtaining (15), we first perform the variation procedure to obtain $p_{m}(\theta)$, then take the limit $m\to\infty$. Switching the order of these two mathematical steps does not lead to the same result. Therefore, K-L divergence is not an appropriate choice.
Appendix D Proof of (36) in Standard Quantum Mechanics